OpenStudy (anonymous):

Find the following: if f(x) = x²+1 and g(x)=2x-3 whats ==> (G*F)(2)

OpenStudy (anonymous):

(gf) (x) = (x^2+1)(2x-3)

OpenStudy (anonymous):

G*F=2(x^2+1)-3

OpenStudy (anonymous):

sub x=2 so 5(1)=5

OpenStudy (anonymous):

put x=2 gives. 7

OpenStudy (anonymous):

(gf)(2) = 5 other guy doesnt know what hes on botu lol

OpenStudy (anonymous):

so its 2(x²+1)-3 => 2x²+2-3 ==>(2x²-1) ===>(2x²-1)(2) =====>(4x²-2)

OpenStudy (anonymous):

elecengineer states that g(f(x)) is simply the producy g(x).f(x), which is incorrect. The composition of two functions g(f(x)) is as stated by dipankarstudy. ghhosst seems to be substituting incorrectly. As he states, after simplifying, g(f(x)) = 2x^2 - 1. Putting x = 2, we get 2(2)^2 - 1 = (2 times 4) - 1 = 7 as derived above by dipankarstudy

OpenStudy (anonymous):

lols whatever guyc newb . I am doind second year engineering noob, i am current on 100% for vector calculus and complex analysis, but you can think what you want

OpenStudy (anonymous):

if they wanted the compositions of function then they would have written g (f(x)) , IN THAT FORM! when someone writes f*g , they have even included the asterick, which denotes multiplication, so if you want to be 100% pedantic ( which I am ) , the asker wants the product of the functions evaluated at 2

OpenStudy (anonymous):

yes, I have seen some people use (g o f ) (x) to represent g(f(x)) , but in this case the asterick was used, suggest a product of functions

OpenStudy (anonymous):

Yes, I see what you mean. It could be that the original poster typed the question slightly differently, thinking that g*f(2) means the same as g(f(2)). Apologies, I didn't mean to be rude.